# Time dependence of neutrino quantum kinetics in a core-collapse supernova

### v1

@@ 5 Discussion
where denotes the energy or or , MeV denotes the -value of the -reaction, MeV is the mass of the electron, is the characteristic neutrino interaction cross section ( cm2), is the electron Pauli-blocking factor, and is the axial coupling.
@@
In order to facilitate the numerical solution of the QKEs, we choose to attenuate the self-interaction strength by a constant factor (see Eq. 2.6), nevertheless for all pot-bounce times we have that at neutrino decoupling. When flavor equipartition is reached, it is unlikely that the rescaling of the self-interaction strength can affect the results. However, for the earlier accretion phase ( s), flavor equipartition is not reached, and it is possible that this scaling might slightly affect the exact final flavor outcome and the time needed to reach a quasi-steady state solution.
@@
We find that flavor equipartition is obtained for s. Such results may seem to confirm recent work in this direction [39, 36, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49]. However, as shown through the flavor configuration obtained by tweaking the collision term in Sec. 4.3, we stress that flavor equipartition depends on the initial conditions and should not be considered a general outcome of global flavor simulations [37, 33, 34].
@@ 6 Conclusions
We adopt the classical steady-state solution to solve the QKEs, this time taking into account neutrino flavor conversion. Fast flavor conversion is responsible for making the spectral energy distribution of () equal to the one of () for s, i.e. flavor equipartition is reached. Despite the changes in the neutrino energy and angular distributions of all flavors, we find that the flavor-dependent neutrino decoupling radii are negligibly affected for the investigated flavor configurations; this is because of the relatively similar properties between the electron and the non-electron flavors characterizing the late accretion phase that facilitate flavor equipartition, but overall negligibly affect the energy and angle integrated flavor-dependent number density at decoupling. On the other hand, ELN crossings are not found for s, but flavor conversion develops because of vacuum mixing, which effectively induces a crossing in the ELN angular distributions.
Remarkably, flavor equipartition is not a generic outcome of fast flavor conversion, but rather the result of the interplay between collisions, flavor conversion and advection, when the neutrino emission properties of the electron and non-electron flavors are reasonably comparable to each other. Modifying the collision term artificially for a time snapshot extracted at s to induce a crossing in the ELN angular distribution, we find that flavor equipartition is not achieved. This is because the differences in the emission properties of different flavors are larger in the early accretion phase and the interplay between flavor conversion and collisions cannot equilibrate the electron and non-electron flavors.
@@ Acknowledgments
This project has received support from the Danmarks Frie Forskningsfond (Project No. 8049-00038B), the European Union (ERC, ANET, Project No. 101087058), and the Deutsche Forschungsgemeinschaft through Sonderforschungbereich SFB 1258 “Neutrinos and Dark Matter in Astro- and Particle Physics” (NDM). Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.

### v2

@@ 5 Discussion
where denotes the energy or or , MeV denotes the -value of the -reaction, MeV is the mass of the electron, is the characteristic neutrino interaction cross section ( cm2), is the electron Pauli-blocking factor, and is the axial coupling. This heating rate has been obtained by integrating the product of cross sections [65] and the energy released per interaction.
@@
In order to facilitate the numerical solution of the QKEs, we choose to attenuate the self-interaction strength by a constant factor (see Eq. 2.6), nevertheless for all post-bounce times we have that at neutrino decoupling. When flavor equipartition is reached, it is unlikely that the rescaling of the self-interaction strength can affect the results. However, for the earlier accretion phase ( s), flavor equipartition is not reached, and it is possible that this scaling might slightly affect the exact final flavor outcome and the time needed to reach a quasi-steady state solution.
@@
We find that flavor equipartition is obtained for s. Such results may seem to confirm recent work in this direction [44, 41, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54], obtained relying on local box simulations and finding near flavor equipartition on one side of the ELN crossing. However, as shown through the flavor configuration obtained by tweaking the collision term in Sec. 4.3, we stress that flavor equipartition depends on the initial conditions and should not be considered a general outcome of global flavor simulations [42, 38, 39].
@@ 6 Conclusions
We adopt the classical steady-state solution to solve the QKEs, this time taking into account neutrino flavor conversion; however, for the sake of simplicity, we neglect the impact of the matter (non-neutrino) background on the flavor evolution. Fast flavor conversion is responsible for making the angular and spectral energy distribution of () equal to the one of () for s, i.e. flavor equipartition is reached. Despite the changes in the neutrino energy and angular distributions of all flavors, we find that the flavor-dependent neutrino decoupling radii are negligibly affected for the investigated flavor configurations; this is because of the relatively similar properties between the electron and non-electron flavors characterizing the late accretion phase that facilitate flavor equipartition, but overall negligibly affect the energy and angle integrated flavor-dependent number density at decoupling. On the other hand, ELN crossings are not found for s, but flavor conversion starts because of slow collective instabilities.
Remarkably, flavor equipartition between the angular and/or energy distributions of the electron and non-electron flavors is not a generic outcome of fast flavor conversion, but rather the result of the interplay between collisions, flavor conversion, and advection, when the neutrino emission properties of the electron and non-electron flavors are reasonably comparable to each other. Modifying the collision term artificially for a time snapshot extracted at s to induce a crossing in the ELN angular distribution, we find that flavor equipartition is not achieved. This is because the differences in the emission properties of different flavors are larger in the early accretion phase and the interplay between flavor conversion and collisions cannot equilibrate the electron and non-electron flavors.
@@ Acknowledgments
We thank Marie Cornelius, Damiano Fiorillo, Georg Raffelt, Guenther Sigl, Meng-Ru Wu, and Zewei Xiong for their comments on the manuscript. This project has received support from the Danmarks Frie Forskningsfond (Project No. 8049-00038B), the European Union (ERC, ANET, Project No. 101087058), and the Deutsche Forschungsgemeinschaft through Sonderforschungbereich SFB 1258 “Neutrinos and Dark Matter in Astro- and Particle Physics” (NDM). Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. The Tycho supercomputer hosted at the SCIENCE HPC Center at the University of Copenhagen was used for supporting the numerical simulations presented in this work.
Appendix A Comparison of the neutrino number densities with the output of the supernova model
In this appendix, we provide a comparison between the number densities resulting from our benchmark hydrodynamical supernova model [55] and the neutrino number densities obtained by solving Eqs. 2.1 and 2.2 with and . Figure 13 shows the angle and energy integrated neutrino number densities for our six selected post-bounce time steps ( and s). The dashed lines show the number densities obtained in this work, whereas the dotted lines show the number densities from the hydrodynamical supernova model [55]. For all times and radii, the agreement is excellent with a difference between the two cases smaller than . We refer the reader to Sec. III and Appendix B of Ref. [26] for more details and an analogous implementation within a single-energy approximation.
We stress, however, that the scope of this work is not to reproduce the outputs of the hydrodynamical simulations with a full Boltzmann approach. We aim to compute the neutrino angular distributions from first principles and explore the variations in the flavor conversion physics.